OpenStudy (anonymous):

what is the arccos(1/sqrt2)

OpenStudy (anonymous):

45

OpenStudy (tkhunny):

\(\cos(What) = \dfrac{1}{\sqrt{2}} = \dfrac{\sqrt{2}}{2}\)?

OpenStudy (tkhunny):

@limkyle Please don't do that. Take some time and read the Code of Conduct.

OpenStudy (anonymous):